# Accelerating Parallel Diffusion Model Serving with Residual Compression

### v1

@@ 1 Introduction
Diffusion models are scaling rapidly both in size and in computation, outpacing the capacity of single accelerators. The model sizes have grown from 983M parameters in Stable Diffusion 1.5 [1] to more than 12B in FLUX.1 [2]. Meanwhile, diffusion models are significantly more computationally intensive than LLMs, with compute increasing faster than the model size [3]. As computational demand rises, single-GPU inference can no longer meet latency constraints, making multiaccelerator parallelism essential for practical deployment.
However, parallel paradigms introduce a new challenge: the increasing prominence of communication bottlenecks, stemming from the fact that interconnect bandwidth has not kept pace with the growth of the compute. From A100 to H100, FP16 FLOPS increases over (312T 1,979T), while NVLink bandwidth grows only and PCIe bandwidth simply doubles [4, 5]. In FLUX.1, standard patch parallelism transmits around 60 GB of activations per image per GPU, consuming over inference time across 4L20 with PCIe interconnects. As bandwidth lags behind, communication increasingly dominates inference cost, underscoring the need to reduce transmission overhead.
Previous works exploited the Temporal Redundancy [6] inherent in diffusion models, where adjacent inference steps produce highly similar activations, to mitigate the communication bottleneck and accelerate parallel inference. Methods like DistriFusion [7] and PipeFusion [8] adopt Displaced Parallelism, which reuses stale activations from previous steps to overlap communication with computation. Although this reduces visible latency, it introduces three fundamental limitations. (1) Displaced parallelism replaces current activations with outdated ones, leading to noticeable quality degradation. (2) Moreover, the core data volume remains unreduced, meaning communication-intensive strategies like sequence parallel [9] still transmit large activations. Consequently, their performance gains are fragile and collapse when the overlap window insufficiently masks communication costs. (3) It requires nontrivial rework of the model pipeline [10], which makes integration complex and limits generality across architectures.
@@
Avoiding stale activations leads to higher quality. Residual Compression works with current data, avoiding the quality degradation inherent in using stale activations in prior methods, yielding significantly higher fidelity. Under aggressive 2-bit quantization, residual compression achieves excellent quality , substantially outperforming state-of-the-art DistriFusion and PipeFusion in identical setups (Figure 1). Remarkably, even at extreme compression ratios of over (transmitting % of the original data), our method maintains better quality than DistriFusion (Figures 2 and 4.2).
A significant drop in data volume yields much lower latency. Residual Compression significantly reduces data exchange, delivering superior end-to-end performance across diverse hardware configurations. It achieved speedup on 4H20 (NVLink) and 4L20 (PCIe) clusters (Figure 1), surpassing prior methods. Furthermore, it makes communication-intensive parallel strategies practical even in low-bandwidth environments, achieving speedup over DistriFusion on Ethernet-level bandwidth.
Residual compression is structurally decoupled from the parallel pipeline, enabling broad compatibility and easy adoption. It operates entirely at the communication layer, without modifying model logic or parallel execution flow. The design generalizes across compression methods (Figure 2) and parallel strategies, and has been applied on large-scale image and video models such as FLUX.1 and CogVideoX, and integrated into frameworks including xDiT and distrifuser, with fewer than 20 lines of core code changed. In contrast, prior methods are tightly coupled with specific strategies and require substantial rework. By focusing exclusively on the transmission aspect, residual compression maintains a lightweight profile and exhibits deployment readiness.
CompactFusion implements residual compression as a portable system, supporting compression techniques including quantization [11], low-rank, and sparsity (as shown in Figure 2), with tunable ratios from to over while preserving high fidelity. It works with parallel strategies such as Patch Parallel and Ring Attention, and integrates into existing frameworks (xDiT, distrifuser) with minimal modification. We believe CompactFusion represents a promising paradigm shift in parallel diffusion acceleration.
@@ 2 Preliminaries on Parallelism
Parallelism is the art of distributing tasks across multiple devices, enabling accelerators to work collaboratively for improved latency and throughput. This distribution can occur across various dimensions: input token(patch) sequence(Sequence Parallel [9]), model layers (Pipeline Parallel [12]), intra-layer computation (Tensor/Expert Parallel [13, 14]), or data samples (Data Parallel [15]). Notably, many parallel strategies necessitate intensive inter-device communication [7] due to the partitioning of weights, activations, or inputs among devices, requiring data exchange for complete computation.
This work focuses on Sequence Parallel [9, 16, 17], which partitions the input sequence across devices and is used in recent works on parallel diffusion [7, 8] as well as in high-performance diffusion inference frameworks [10, 18]. It is by far the most widely adopted parallel strategy for diffusion models, as it is highly latency-friendly and well suited for real-time generation.
@@ 3 Method
CompactFusion mitigates the communication bottleneck in parallel diffusion by leveraging residual compression (Sections 3.1, 3.2 and 3.3) and practical system co-design (Section 3.4), achieving significant data reduction with minimal quality degradation.
@@ 3.1 Exploiting Temporal Redundancy via Residual Compression
Transmitting compressed residuals effectively eliminates redundancy from the data, lowering the communication volume while preserving fidelity. This efficiency stems from a key insight: diffusion models exhibit high temporal redundancy, with activations changing only minimally between steps. As a result, transmitting full activations often involves sending largely redundant information. By encoding only the meaningful differences, we can greatly reduce the volume of transmitted data. However, naively compressing high-dimensional activations still introduces substantial errors, often resulting in severe quality degradation or visual collapse (as shown in Figure 4(a)). Effectively leveraging temporal redundancy, therefore, requires more than simply reducing the amount of data—it demands meticulous design.
Our optimization builds on two observations. First, activations change gradually across steps, their step-wise residuals(differences from the previous result) having much smaller magnitude (as shown in Figure 5, residual norm is only a fraction of full activation norm). Second, compressing a smaller-magnitude signal typically induces proportionally less error. This makes the residual a more efficient compression target as it captures the essential change. We therefore propose residual compression: instead of transmitting the full activation, we transmit its compressed difference from the previous reconstruction and recover the current state by adding it back.
To verify this principle, we apply aggressive 1-bit binarization(detailed in Appendix C) to both naive (compressing full activation, Figure 4(b)) and residual compression (Figure 4(c)) over Ring Attention. Naive compression fails completely: the output is corrupted, with large errors. In contrast, residual compression yields clean reconstructions with much smaller error at each step, and the activation trajectory remains close to the uncompressed baseline throughout.
While the idea of transmitting differences exists in other contexts [20, 21], to the best of our knowledge, CompactFusion is the first to demonstrate its effectiveness on parallel inference, and also the first on diffusion models, enabling aggressive data reduction while preserving generation quality.
@@ 3.2 Compensating Error Accumulation with Feedback
Residual compression significantly reduces per-step compression error by targeting smaller-magnitude signals. However, these small errors still accumulate over diffusion steps (curve for Total Error, Figure 4(c)). To eliminate long-term drift and preserve fidelity, CompactFusion integrates Error Feedback, a technique used in gradient compression [22, 23]. Rather than discarding the residual compression error at each step, we store it locally and add it to the next-step residual before compression, recirculating untransmitted information and systematically compensating for loss (illustrated in Figure 3(a)). This stabilizes reconstruction and prevents the compressed state from diverging from the original trajectory. Details are in Appendix G.
Problem Formulation. To understand how compression affects inference quality, we model diffusion as a sequence of transformations , where denotes the intermediate activation in a fixed layer and captures the combined effect of one forward pass and scheduler update. Following relevant works [24, 25]. we define a -compressor with , satisfying for .
Assumptions. Our analysis relies on realistic assumptions grounded in empirical observations of diffusion behavior. First, the model exhibits strong local stability—captured by -smoothness of each transformation , i.e., , with . Second, activations evolve gradually over time due to temporal redundancy: the expected per-step change is much smaller in scale than the activations themselves, i.e., . We also assume uncorrelated error terms where appropriate. The definitions and assumptions are elaborated in Section G.2 and Section G.3; these properties are well-supported in practice—activation norms remain bounded, and residual magnitude is significantly smaller than that of activations (Figure 5).
@@
See Appendix G for the proof. This ratio confirms that residual compression with feedback yields a significantly lower steady-state error, often by an order of magnitude or more. The result holds under a mild stability condition: which imposes a lower bound on the required compression quality. In contrast, Residual compression without error feedback fails to converge; its error grows linearly with time steps and does not admit a steady state bound.
Validation. We validate this theory empirically using 1-bit binarization. As shown in Figure 4(d), CompactFusion with error feedback maintains low cumulative deviation throughout the inference process, closely follows the original trajectory, and generates high-quality images.
@@ 3.3 Scaling Residual Compression to Extreme Ratios
To make this feasible under real-time constraints, we adopt subspace iteration [26] (detailed in Appendix C), a faster alternative to SVD. While SVD offers optimality, it is approximately times slower than the subspace iteration solution, far exceeding the 5ms communication window and 1ms compression budget in practice. In contrast, we adapt the subspace iteration from gradient compression [27], producing a usable approximation in milliseconds.
@@
We empirically validate this strategy by comparing it with an alternative that improves approximation optimality through more subspace iterations(see Section E.1). Despite added quantization error, expanding rank coverage yields significantly better generation quality, confirming that directional breadth matters more than per-step accuracy under tight bandwidth constraints.
@@ 3.4 System Co-Design for Efficiency and Usability
We develop highly optimized GPU kernels for our compression techniques. Beyond efficient quantization and subspace iteration, we repurpose the concept of N:M sparsity [28] which originally used in weight pruning, and adapt it to activation compression via a N:M block sparsifier. Unlike standard TopK methods that suffer from sorting overhead and irregular memory access, our block sparsifier enables linear-time selection and GPU-friendly memory patterns. It is the only sparsifier design we found to yield practical on-device speedup.
@@
CompactFusion benefits from the inherent modularity of Residual Compression, which cleanly separates compression from execution logic (Figure 3(b)). It wraps standard communication primitives without altering model code or parallel pipeline, and generalizes across strategies, compressors, and frameworks. In contrast, prior methods often entangle with specific parallel strategies and require substantial pipeline rework.
@@ 4.1 Setups
To demonstrate broad applicability, experiments are conducted on various hardware and interconnects: high-bandwidth NVLink (H20 clusters, bandwidth: 366 GB/s), standard PCIe (L20 clusters, bandwidth: 17.13 GB/s) and simulated lower-bandwidth Ethernet (A40 clusters using tc for trafic control), ensuring robustness evaluation under various deployment constraints.
@@
We apply CompactFusion to Sequence Parallelism, the most widely adopted strategy in parallel diffusion (Section 2), to reflect real-world deployment scenarios. We compare it with strong baselines from both industry and research, including Patch Parallel (all-gather KV) [18], Ring Attention (ring-style KV transfer) [9], DeepSpeed-Ulysses (all-to-all) [16], and DistriFusion (displaced patch parallel) [18]. We also include PipeFusion (displaced TeraPipe [32]) [8] as a baseline, though it is orthogonal to our method. We note that PipeFusion is not evaluated on video models as the xDiT does not support PipeFusion for video models (further discussed in Appendix B).
@@
Following the relevant work, we adopt the standard evaluataion metrics for image quality testing: Peak Signal Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [33] and Fréchet Inception Distance (FID) [34]. We adopt the following metrics for video quality testing: Structural Similarity Index Measure (SSIM)[35], Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS)[33]. We also conduct a human evaluation study to assess perceptual consistency from a human perspective (detailed in Section D.1).
@@ 4.2 Main Results
CompactFusion achieves lower latency while maintaining high visual fidelity. It consistently performs well across different generation models, including FLUX.1-dev for images and CogVideoX for videos. It is also robust across hardware setups—L20, H20, and A40—and across network conditions such as NVLink, PCIe, and simulated Ethernet. These results are shown in Figure 6. Quantitative metrics across 3, 4, and 6-GPU scales are provided in Tables 1 and 2.
Notably, CompactFusion uniquely enables the widely adopted but communication-intensive sequence parallel over slow networks, achieving up to speedup over DistriFusion under Ethernet-level bandwidth conditions (Figure 7).
By directly eliminating redundant data, CompactFusion drastically reduces the transmission volume. In its low-rank variant, it sends less than 1% of the original activation data. Despite this, it still outperforms DistriFusion in generation quality, as shown in Table 1 (image) and Table 2 (video).

### v2

@@ 1 Introduction
Diffusion models are scaling rapidly both in size and in computation, outpacing the capacity of single accelerators. The model sizes have grown from 983M parameters in Stable Diffusion 1.5 [sd15] to more than 12B in FLUX.1 [flux]. Meanwhile, diffusion models are significantly more computationally intensive than LLMs, with compute increasing faster than the model size [svdquant]. As computational demand rises, single-GPU inference can no longer meet latency constraints, making multiaccelerator parallelism essential for practical deployment.
However, parallel paradigms introduce a new challenge: the increasing prominence of communication bottlenecks, stemming from the fact that interconnect bandwidth has not kept pace with the growth of the compute. From A100 to H100, FP16 FLOPS increases over (312T 1,979T), while NVLink bandwidth grows only and PCIe bandwidth simply doubles [a100, h100]. In FLUX.1, standard patch parallelism transmits around 60 GB of activations per image per GPU, consuming over inference time across 4L20 with PCIe interconnects. As bandwidth lags behind, communication increasingly dominates inference cost, underscoring the need to reduce transmission overhead.
Previous works exploited the Temporal Redundancy [redundancy] inherent in diffusion models, where adjacent inference steps produce highly similar activations, to mitigate the communication bottleneck and accelerate parallel inference. Methods like DistriFusion [distrifusion] and PipeFusion [pipefusion] adopt Displaced Parallelism, which reuses stale activations from previous steps to overlap communication with computation. Although this reduces visible latency, it introduces three fundamental limitations. (1) Displaced parallelism replaces current activations with outdated ones, leading to noticeable quality degradation. (2) Moreover, the core data volume remains unreduced, meaning communication-intensive strategies like sequence parallel [ring] still transmit large activations. Consequently, their performance gains are fragile and collapse when the overlap window insufficiently masks communication costs. (3) It requires nontrivial rework of the model pipeline [xfuser], which makes integration complex and limits generality across architectures.
@@
Avoiding stale activations leads to higher quality. Residual Compression works with current data, avoiding the quality degradation inherent in using stale activations in prior methods, yielding significantly higher fidelity. Under aggressive 2-bit quantization, residual compression achieves excellent quality , substantially outperforming state-of-the-art DistriFusion and PipeFusion in identical setups (Figure˜1). Remarkably, even at extreme compression ratios of over (transmitting % of the original data), our method maintains better quality than DistriFusion (Figures˜2 and 4.2).
A significant drop in data volume yields much lower latency. Residual Compression significantly reduces data exchange, delivering superior end-to-end performance across diverse hardware configurations. It achieved speedup on 4H20 (NVLink) and 4L20 (PCIe) clusters (Figure˜1), surpassing prior methods. Furthermore, it makes communication-intensive parallel strategies practical even in low-bandwidth environments, achieving speedup over DistriFusion on Ethernet-level bandwidth.
Residual compression is structurally decoupled from the parallel pipeline, enabling broad compatibility and easy adoption. It operates entirely at the communication layer, without modifying model logic or parallel execution flow. The design generalizes across compression methods (Figure˜2) and parallel strategies, and has been applied on large-scale image and video models such as FLUX.1 and CogVideoX, and integrated into frameworks including xDiT and distrifuser, with fewer than 20 lines of core code changed. In contrast, prior methods are tightly coupled with specific strategies and require substantial rework. By focusing exclusively on the transmission aspect, residual compression maintains a lightweight profile and exhibits deployment readiness.
CompactFusion implements residual compression as a portable system, supporting compression techniques including quantization [quant], low-rank, and sparsity (as shown in Figure˜2), with tunable ratios from to over while preserving high fidelity. It works with parallel strategies such as Patch Parallel and Ring Attention, and integrates into existing frameworks (xDiT, distrifuser) with minimal modification. We believe CompactFusion represents a promising paradigm shift in parallel diffusion acceleration.
@@ 2 Preliminaries on Parallelism
Parallelism is the art of distributing tasks across multiple devices, enabling accelerators to work collaboratively for improved latency and throughput. This distribution can occur across various dimensions: input token(patch) sequence(Sequence Parallel [ring]), model layers (Pipeline Parallel [gpipe]), intra-layer computation (Tensor/Expert Parallel [megatron, gshard]), or data samples (Data Parallel [zero]). Notably, many parallel strategies necessitate intensive inter-device communication [distrifusion] due to the partitioning of weights, activations, or inputs among devices, requiring data exchange for complete computation.
This work focuses on Sequence Parallel [ring, ulysses, usp], which partitions the input sequence across devices and is used in recent works on parallel diffusion [distrifusion, pipefusion] as well as in high-performance diffusion inference frameworks [xfuser, distrifuser]. It is by far the most widely adopted parallel strategy for diffusion models, as it is highly latency-friendly and well suited for real-time generation. A more detailed description of these strategies is provided in Appendix˜B.
To reduce perceived latency, prior works overlaps communication with computation by reusing stale activations [distrifusion, pipefusion, pcpp], but suffers from quality degradation, complex pipeline rework, and unchanged (potentially large) communication volume, limiting its effectiveness.
@@ 3 Method
CompactFusion mitigates the communication bottleneck in parallel diffusion by leveraging residual compression (Sections˜3.1, 3.2 and 3.3) and practical system co-design (Section˜3.4), achieving significant data reduction with minimal quality degradation.
@@ 3.1 Exploiting Temporal Redundancy via Residual Compression
Transmitting compressed residuals effectively eliminates redundancy from the data, lowering the communication volume while preserving fidelity. This efficiency stems from a key insight: diffusion models exhibit high temporal redundancy, with activations changing only minimally between steps. As a result, transmitting full activations often involves sending largely redundant information. By encoding only the meaningful differences, we can greatly reduce the volume of transmitted data. However, naively compressing high-dimensional activations still introduces substantial errors, often resulting in severe quality degradation or visual collapse (as shown in Figure˜4(a)). Effectively leveraging temporal redundancy, therefore, requires more than simply reducing the amount of data—it demands meticulous design.
Our optimization builds on two observations. First, activations change gradually across steps, their step-wise residuals(differences from the previous result) having much smaller magnitude (as shown in Figure˜5, residual norm is only a fraction of full activation norm). Second, compressing a smaller-magnitude signal typically induces proportionally less error. This makes the residual a more efficient compression target as it captures the essential change. We therefore propose residual compression: instead of transmitting the full activation, we transmit its compressed difference from the previous reconstruction and recover the current state by adding it back.
To verify this principle, we apply aggressive 1-bit binarization (detailed in Appendix˜D) to both naive (compressing full activation, Figure˜4(b)) and residual compression (Figure˜4(c)) over Ring Attention. Naive compression fails completely: the output is corrupted, with large errors. In contrast, residual compression yields clean reconstructions with much smaller error at each step, and the activation trajectory remains close to the uncompressed baseline throughout.
While the idea of transmitting differences exists in other contexts [aqsgd, cachegen], to the best of our knowledge, CompactFusion is the first to demonstrate its effectiveness on parallel inference, and also the first on diffusion models, enabling aggressive data reduction while preserving generation quality.
@@ 3.2 Compensating Error Accumulation with Feedback
Residual compression significantly reduces per-step compression error by targeting smaller-magnitude signals. However, these small errors still accumulate over diffusion steps (curve for Total Error, Figure˜4(c)). To eliminate long-term drift and preserve fidelity, CompactFusion integrates Error Feedback, a technique used in gradient compression [feedback, deepgradcomp]. Rather than discarding the residual compression error at each step, we store it locally and add it to the next-step residual before compression, recirculating untransmitted information and systematically compensating for loss (illustrated in Figure˜3(a)). This stabilizes reconstruction and prevents the compressed state from diverging from the original trajectory. Details are in Appendix˜H.
Problem Formulation. To understand how compression affects inference quality, we model diffusion as a sequence of transformations , where denotes the intermediate activation in a fixed layer and captures the combined effect of one forward pass and scheduler update. Following relevant works [commcompgrad, dagc]. we define a -compressor with , satisfying for .
Assumptions. Our analysis relies on realistic assumptions grounded in empirical observations of diffusion behavior. First, the model exhibits strong local stability—captured by -smoothness of each transformation , i.e., , with . Second, activations evolve gradually over time due to temporal redundancy: the expected per-step change is much smaller in scale than the activations themselves, i.e., . We also assume uncorrelated error terms where appropriate. The definitions and assumptions are elaborated in Section˜H.2 and Section˜H.3; these properties are well-supported in practice—activation norms remain bounded, and residual magnitude is significantly smaller than that of activations (Figure˜5).
@@
See Appendix˜H for the proof. This ratio confirms that residual compression with feedback yields a significantly lower steady-state error, often by an order of magnitude or more. The result holds under a mild stability condition: which imposes a lower bound on the required compression quality. In contrast, Residual compression without error feedback fails to converge; its error grows linearly with time steps and does not admit a steady state bound.
Validation. We validate this theory empirically using 1-bit binarization. As shown in Figure˜4(d), CompactFusion with error feedback maintains low cumulative deviation throughout the inference process, closely follows the original trajectory, and generates high-quality images.
@@ 3.3 Scaling Residual Compression to Extreme Ratios
To make this feasible under real-time constraints, we adopt subspace iteration [subspace] (detailed in Appendix˜D), a faster alternative to SVD. While SVD offers optimality, it is approximately times slower than the subspace iteration solution, far exceeding the 5ms communication window and 1ms compression budget in practice. In contrast, we adapt the subspace iteration from gradient compression [powersgd], producing a usable approximation in milliseconds.
@@
We empirically validate this strategy by comparing it with an alternative that improves approximation optimality through more subspace iterations(see Section˜F.1). Despite added quantization error, expanding rank coverage yields significantly better generation quality, confirming that directional breadth matters more than per-step accuracy under tight bandwidth constraints.
@@ 3.4 System Co-Design for Efficiency and Usability
We develop highly optimized GPU kernels for our compression techniques. Beyond efficient quantization and subspace iteration, we repurpose the concept of N:M sparsity [nmsparse] which originally used in weight pruning, and adapt it to activation compression via a N:M block sparsifier. Unlike standard TopK methods that suffer from sorting overhead and irregular memory access, our block sparsifier enables linear-time selection and GPU-friendly memory patterns. It is the only sparsifier design we found to yield practical on-device speedup.
@@
CompactFusion benefits from the inherent modularity of Residual Compression, which cleanly separates compression from execution logic (Figure˜3(b)). It wraps standard communication primitives without altering model code or parallel pipeline, and generalizes across strategies, compressors, and frameworks. In contrast, prior methods often entangle with specific parallel strategies and require substantial pipeline rework.
@@ 4.1 Setups
Our method works with off-the-shelf models, we evaluate it on the state-of-the-art FLUX.1-dev [flux] for image generation, and on CogVideoX-2b [cogvideo] for video generation. We adhere to standard inference settings on xDiT, employing 28 steps for FLUX.1-dev and 50 steps for CogVideoX-2b.
@@
We test the image generation model using prompts in COCO Captions 2014 dataset [cococaption] and video generation using prompts sampled from VBench [vbench].
@@
To demonstrate broad applicability, experiments are conducted on various hardware and interconnects: high-bandwidth NVLink (H20 clusters, bandwidth: 366 GB/s), standard PCIe (L20 clusters, bandwidth: 17.13 GB/s) and simulated lower-bandwidth Ethernet (A40 clusters using tc for traffic control), ensuring robustness evaluation under various deployment constraints.
@@
We apply CompactFusion to Sequence Parallelism, the most widely adopted strategy in parallel diffusion (Section˜2), to reflect real-world deployment scenarios. We compare it with strong baselines from both industry and research, including Patch Parallel (all-gather KV) [distrifuser], Ring Attention (ring-style KV transfer) [ring], DeepSpeed-Ulysses (all-to-all) [ulysses], and DistriFusion (displaced patch parallel) [distrifuser]. We also include PipeFusion (displaced TeraPipe [terapipe]) [pipefusion] as a baseline, though it is orthogonal to our method. We note that PipeFusion is not evaluated on video models as the xDiT does not support PipeFusion for video models (further discussed in Appendix˜C).
@@
Following the relevant work, we adopt the standard evaluataion metrics for image quality testing: Peak Signal Noise Ratio (PSNR), Learned Perceptual Image Patch Similarity (LPIPS) [lpips] and Fréchet Inception Distance (FID) [fid]. We adopt the following metrics for video quality testing: Structural Similarity Index Measure (SSIM)[ssim], Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS)[lpips]. We also conduct a human evaluation study to assess perceptual consistency from a human perspective (detailed in Section˜E.1).
@@ 4.2 Main Results
CompactFusion achieves lower latency while maintaining high visual fidelity. It consistently performs well across different generation models, including FLUX.1-dev for images and CogVideoX for videos. It is also robust across hardware setups—L20, H20, and A40—and across network conditions such as NVLink, PCIe, and simulated Ethernet. These results are shown in Figure˜6. Quantitative metrics across 3, 4, and 6-GPU scales are provided in Tables˜1 and 2.
Notably, CompactFusion uniquely enables the widely adopted but communication-intensive sequence parallel over slow networks, achieving up to speedup over DistriFusion under Ethernet-level bandwidth conditions (Figure˜7).
By directly eliminating redundant data, CompactFusion drastically reduces the transmission volume. In its low-rank variant, it sends less than 1% of the original activation data. Despite this, it still outperforms DistriFusion in generation quality, as shown in Table˜1 (image) and Table˜2 (video).